# Maximum entropy distributions of wavefunctions at thermal equilibrium

### 19 Mar 2026 v1

@@ Abstract
Statistical mechanics reveals that the properties of a macroscopic physical system emerge as an average over an ensemble of statistically independent microscopic subsystems, each occupying a specific microstate. In some models of quantum systems, these microstates are the wavefunction states of individual quantum systems. The physical principles that govern the distribution of a wavefunction ensemble, even under conditions of thermal equilibrium, are not well established. For instance, the canonical Boltzmann distribution cannot be applied to wavefunctions because they lack a definite energy. In this manuscript, we present a maximum entropy principle for the quantum wavefunction ensemble at thermal equilibrium, the so-called Scrooge ensemble. We highlight that a constraint on the energy expectation value, or even the shape of the associated eigenstate distribution, fails to yield a valid equilibrium state. We find that in addition to these constraints, one must also constrain the measurement entropy to be equal to the Rényi divergence of the ensemble with respect to the Gibbs state, indicating that the Rényi divergence may have uninvestigated physical importance to thermal equilibrium in quantum systems.
@@ Introduction
The statistical mechanical formulation of thermal equilibrium is founded on the maximum entropy principle. This principle states that the equilibrium distribution of system microstates is the one that maximizes entropy subject to the system’s physical constraints.[26, 25] For quantum systems, these microstates are traditionally taken to be the set of energy eigenstates. When the average energy of the eigenstate ensemble is constrained, entropy maximization yields the well-known Gibbs state, in which eigenstates are Boltzmann distributed, as represented by the density matrix,
@@
It is worthwhile to note that there has been skepticism surrounding the seemingly straightforward extension of statistical mechanical principles to wavefunction ensembles. Most notably, it was argued by Jaynes that a distribution of pure state wavefunctions, cannot serve as a valid foundation for statistical mechanics because the state distribution function, $P(\Gamma)$, does not define the probabilities of mutually exclusive events.[25] Jaynes argued that wavefunction states are only mutually exclusive if they are mutually orthogonal, and therefore, a quantum state probability distribution is only valid if it is restricted to an orthonormal set of states. We argue, however, that although non-orthogonal wavefunction states are observationally indistinguishable, they are formally distinct quantum states. For example, they are precisely distinguishable (and therefore mutually exclusive) in the simulation of model quantum systems. If a statistical mechanical formulation of a simulated ensemble of pure state wavefunctions is possible, then it is, at least in principle, physically meaningful and worthy of scientific study (even if these studies are purely theoretical).
@@
Additional information about the ensemble is contained within the higher moments of $P(\Gamma)$. This information is, in principle, experimentally accessible, for example, with experimental protocols that utilize quantum information platforms,[15, 23, 31, 38] or single-molecule spectroscopy.[17, 18, 16]
The mapping from $P(\Gamma)$ to $\rho$ is not one-to-one, i.e., there are many possible wavefunction ensembles capable of representing a given ensemble density matrix, $\rho$. The first attempt to derive an expression for $P(\Gamma)$ was by Park and Band, who defined an entropy similar to eq.˜3, but defined over all possible density matrices, which they took as the fundamental physical states of quantum systems.[35, 6, 36, 7] Their formalism was later refined by Slater, who reduced the probability space to only include pure states;[42, 44] however, his analysis was restricted to small spin systems.[43, 45] Later, Brody and Hughston considered the full $P$-ensemble, maximizing $S_{P}$ subject to a constraint on the average energy expectation value. In addition, they considered the maximum $S_{P}$ distribution for a two-level system subject to a constraint to the Gibbs state.[10, 11, 12, 13, 9] Generalizations to $N$-level systems were made soon after for the energy-constrained distribution,[8, 34] and more recently for the Gibbs state-constrained distribution.[4] [^1] We catch a minor error in the previously published GCE distribution by providing an alternative proof in section S1.2 of the SI.
Here, we critically evaluate and build upon these previous efforts with the goal of establishing a maximum entropy principle for the equilibrium wavefunction ensemble. Before proceeding, we define our criteria for equilibrium. We will consider $P(\Gamma)$ to be a proper thermal distribution if it satisfies the three criteria originally proposed by Goldstein et al.[20] These criteria for equilibrium are:
@@
    $P(\Gamma)$ obeys the hereditary property. That is, if a total system composed of two independent subsystems (e.g. a system and a bath) has its bath projected out (e.g., by performing a measurement on the bath), then the projected system is time invariant and in the Gibbs state with the same distributional form. The hereditary property ensures that the thermal distribution is inherited by the system when measurements occur in the bath of a system-bath composite.
@@
Figure 1: 345.0pt433.62ptThe maximum entropy distributions considered as candidates for the thermal distribution of wavefunctions. (a) An ensemble of wavefunctions $P(\Gamma)$ can be compactly represented by a density operator $\rho$. Maximizing the entropy of $P(\Gamma)$ is not equivalent to maximizing the entropy of $\rho$, subject to the same constraint. (b) The energy constrained ensemble (ECE) is the quantum analogue of the classical canonical ensemble, maximizing entropy under an average energy constraint. (c) The Gibbs-state constrained ensemble (GCE) maximizes entropy among distributions that average to the Gibbs state. (d) The Scrooge ensemble maximizes entropy under a constraint on the average Rényi divergence $D(\Gamma\|\rho)$ of $\Gamma$ from $\rho$.
@@
The Scrooge ensemble, is given by the following expression:
$P_{\mathrm{Scr}}(\Gamma)=\frac{1}{(2\pi)^{N-1}}\frac{N!}{\det\rho}\left(\mathrm{Tr}\{\Gamma\rho^{-1}\}\right)^{-(N+1)}.$ (5)
This distribution was first derived by Jozsa et al. as the solution to Eq. 4 that minimizes accessible information about the system’s wavefunction, thereby being the most ‘stingy’ with information.[27] Later, $P_{\mathrm{Scr}}(\Gamma)$ was independently derived by Goldstein et al., who demonstrated that when $\rho=\rho_{\mathrm{G}}$, $P_{\mathrm{Scr}}(\Gamma)$ uniquely satisfies the above three criteria for proper thermal equilibrium.[20] Previous work has demonstrated that the Scrooge ensemble has maximum entropy in a certain sense, as it can be obtained from a maximum entropy distribution over unnormalized quantum states (subject to a density matrix constraint), by projecting the unnormalized states to the unit sphere.[20, 31] However, the physical interpretation of this demonstration is not clear. If the Scrooge ensemble can be formulated from a maximum entropy principle on normalized wavefunctions, then the associated constraints may guide our understanding of its physical origins.
As we now demonstrate, the Scrooge ensemble is the maximum entropy distribution subject to a constraint on the Rényi divergence. The Rényi divergence is a tunable measure of the difference between two probability distributions or density matrices.[39, 37] Specifically, the Rényi divergence between a pure state density matrix $\Gamma$, and an ensemble density matrix, $\rho$, is given by,
@@
where $\alpha$ is a tunable parameter capable of shifting the emphasis of the divergence between the features of $\Gamma$ and $\rho$. For $\alpha\rightarrow 0$, Eq. 6 simply measures the overlap between $\Gamma$ and $\rho$. In the limit that $\alpha\rightarrow 1$, the Rényi divergence reduces to the Kullback–Leibler divergence, or equivalently to Umegaki’s relative entropy[46] ($\lim_{\alpha\rightarrow 1}D_{\alpha}(\Gamma\|\rho)=-\mathrm{Tr}\{\Gamma\ln\rho\}$). As we increase $\alpha\rightarrow\infty$, the divergence becomes more stringent and sensitive towards the largest population difference between $\Gamma$ and $\rho$.[47]
@@
where $C(\rho)$ is an unspecified constraint function. We then determine the specific form of $C(\rho)$ that yields the equilibrium Scrooge ensemble. Using the method of Lagrange multipliers, the maximum entropy distribution is given by (proof in section S1.3 of the SI),
@@
where $\mu$ is the Lagrange multiplier enforcing the constraint $\langle D_{\alpha}(\Gamma\|\rho)\rangle_{P_{\alpha\mu}}=C(\rho)$. Comparing Eq. 8 to Eq. 5 reveals that this maximum entropy solution yields the Scrooge ensemble in the case where $\alpha=2$ and $\mu=N+1$. Substituting these parameter values into $P_{\alpha\mu}(\Gamma)$, and using the resulting distribution to compute the average in Eq. 7 provides a solution to the constraint (proof in section S1.3 of the SI),
@@
In section S2 of the SI, we present a proof that the constraint in Eq. 13 is the unique constraint on the Rényi divergence that is self-consistent, elevating the Scrooge ensemble as an important maximum entropy solution among this class of constraints. That is, this particular constraint applied specifically with $\alpha=2$ is the only Rényi divergence-based constraint with a maximum entropy principle corresponding to thermal equilibrium. A consequence of this proof is that it demonstrates that any constrained Rényi divergence between the density operator and its composing wavefunctions must yield the Scrooge ensemble in a maximum entropy procedure. However, it is possible that different physical constraints exist that equivalently yield the Scrooge ensemble as a maximum entropy solution.
The constraint in Eq. 13 includes a hierarchy of indirect constraints. This includes a constraint that the first moment of $P(\Gamma)$ is equal to the Gibbs state, $\rho_{\mathrm{G}}$, which, in turn, includes an indirect constraint on the average energy of the wavefunction ensemble. In the following sections, we descend this hierarchy of constraints and evaluate the consequences for the associated maximum entropy principles.
@@
where $\bar{H}(\Gamma)=\mathrm{Tr}\{H\Gamma\}$ is the energy of state $\Gamma$, and $\beta^{\prime}$ is a statistical inverse temperature that depends on the average energy. Notably, $\beta^{\prime}$ differs from the thermodynamic inverse temperature, $\beta$. However, the two quantities can be related through the energy constraint expression (proof in section S1.1 of the SI),[22, 24, 32]
$\frac{\sum_{i}E_{i}\alpha_{i}}{\sum_{i}\alpha_{i}}+\frac{N-1}{\beta^{\prime}}=\frac{\sum_{i}E_{i}e^{-\beta E_{i}}}{\sum_{i}e^{-\beta E_{i}}},\quad\alpha_{i}=\frac{e^{-\beta^{\prime}E_{i}}}{\prod_{j\neq i}(\beta^{\prime}E_{j}-\beta^{\prime}E_{i})}$ (15)
where $E_{i}$ denotes the eigenvalue of the $i$th energy eigenstate.
@@
Figure 2: The populations of the ECE (red stars) do not agree with the Gibbs state (black stars). (a) The thermal populations at high temperature are similar. (b) At low temperature, the ground state and higher excited state populations are increased in the ECE. Inset shows populations on a logarithmic scale. Monte Carlo sampling of the thermal equilibrium distribution (pink circles) agrees with the analytic results. Results were calculated for the Hamiltonian spectrum $\sigma\{H\}=\{k\}_{k=0}^{9}$.
@@
Although the Gibbs state is preserved, we demonstrate that the GCE violates the hereditary property and thus cannot be a valid thermal distribution. The hereditary property requires the projected ensemble on the system (in a system-bath composite) to be the same as the thermal distribution on the system.[20] More concretely, let $|\Psi\rangle\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ be a wavefunction sampled from the thermal distribution $P(\Gamma)$ on the composite Hilbert space of $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$. Then the projected ensemble on $\mathcal{H}_{1}$ given a measurement basis $\{|q_{2}\rangle\}$ for $\mathcal{H}_{2}$ is
@@
Figure 3: The hereditary property is violated by the GCE and satisfied by the Scrooge ensemble. The Kolmogov-Smirnov (KS) statistic of the projected ensemble and thermal distribution of the system as a function of inverse temperature $\beta$. The GCE satisfies the KS test at high temperatures for the significance levels of $\alpha=5\%$ and $0.5\%$, but fails at low temperatures. The Scrooge ensemble satisfies the hereditary property for all temperatures.[20] Results are calculated by Monte Carlo sampling, with the system and bath spectra given by $\sigma\{H\}=\{0,1\}$. Error bars are standard error in the mean over 1010 trials with $10^{7}$ samples.
@@ Discussion
The maximum entropy constraint of the Scrooge ensemble currently lacks a strong a priori physical justification. Due to the uniqueness of the constraint $\langle D_{2}(\Gamma\|\rho)\rangle_{P}=\langle H_{A}(\rho)\rangle_{A}$ amongst the class of divergence constraints, justifying it reduces to rationalizing why the average excess surprisal, or loosely speaking, the ‘distance’ of wavefunctions from their average $\rho$, should depend on $\rho$ at thermal equilibrium.

### 20 Mar 2026 v2

@@ Abstract
Statistical mechanics reveals that the properties of a macroscopic physical system emerge as an average over an ensemble of statistically independent microscopic subsystems, each occupying a specific microstate. In the study of quantum systems, these microstates can be chosen to correspond to the pure state wavefunctions of individual quantum systems. However, the physical principles that govern the distribution of a pure state wavefunction ensemble, even under conditions of thermal equilibrium, are not well established. For instance, the canonical Boltzmann distribution cannot be applied to wavefunctions because they lack a definite energy. In this manuscript, we present a maximum entropy principle for the quantum wavefunction ensemble at thermal equilibrium, the so-called Scrooge ensemble. We highlight that a constraint on the energy expectation value, or even the shape of the associated eigenstate distribution, fails to yield a valid equilibrium state. We find that in addition to these constraints, one must also constrain the measurement entropy to be equal to the Rényi divergence of the ensemble with respect to the Gibbs state, indicating that the Rényi divergence may have uninvestigated physical importance to thermal equilibrium in quantum systems.
@@ Introduction
The statistical mechanical formulation of thermal equilibrium is founded on the maximum entropy principle. This principle states that the equilibrium distribution of system microstates is the one that maximizes entropy subject to the system’s physical constraints.[26, 25] For quantum systems represented in a basis of eigenstates, when the average energy of the eigenstate ensemble is constrained, entropy maximization yields the well-known Gibbs state, in which eigenstates are Boltzmann distributed, as represented by the density matrix,
@@
It is worthwhile to note that there has been skepticism surrounding the seemingly straightforward extension of statistical mechanical principles to wavefunction ensembles. Most notably, it was argued by Jaynes that a distribution of pure state wavefunctions, cannot serve as a valid foundation for statistical mechanics because the state distribution function, $P(\Gamma)$, does not define the probabilities of mutually exclusive events.[25] Jaynes argued that wavefunction states are only mutually exclusive if they are orthogonal, and therefore, a quantum state probability distribution is only valid if it is restricted to an orthonormal set of states. We argue, however, that although non-orthogonal wavefunction states are observationally indistinguishable, they are formally distinct quantum states. For example, they are precisely distinguishable (and therefore mutually exclusive) in the numerical simulation of model quantum systems. If a statistical mechanical formulation of a simulated ensemble of pure state wavefunctions is possible, then it is, at least in principle, physically meaningful and worthy of study (even if these studies are purely theoretical).
Figure 1: Maximum entropy distribution derived from different physical constraints. (a) An ensemble of wavefunctions $P(\Gamma)$ can be compactly represented by a density operator $\rho$. Maximizing the entropy of $P(\Gamma)$ is not equivalent to maximizing the entropy of $\rho$, subject to the same constraint. (b) The energy constrained ensemble (ECE) is the quantum analogue of the classical canonical ensemble, maximizing entropy under an average energy constraint. (c) The Gibbs-state constrained ensemble (GCE) maximizes entropy among distributions that average to the Gibbs state. (d) The Scrooge ensemble maximizes entropy under a constraint on the average Rényi divergence $D(\Gamma\|\rho)$ of $\Gamma$ from $\rho$.
@@
Furthermore, additional information about the ensemble is contained within the higher moments of $P(\Gamma)$. This information is, in principle, experimentally accessible, for example, with experimental protocols that utilize quantum information platforms,[15, 23, 31, 38] or single-molecule spectroscopy.[17, 18, 16]
The mapping from $P(\Gamma)$ to $\rho$ is not one-to-one, i.e., there are many possible wavefunction ensembles capable of representing a given ensemble density matrix, $\rho$. The first attempt to derive an expression for $P(\Gamma)$ at equilibrium was by Park and Band, who defined an entropy similar to eq.˜3, but defined over all possible density matrices, which they took as the fundamental physical states of quantum systems.[35, 6, 36, 7] Their formalism was later refined by Slater, who reduced the probability space to only include pure states;[42, 44] however, his analysis was restricted to small spin systems.[43, 45] Later, Brody and Hughston considered the full $P$-ensemble, maximizing $S_{P}$ subject to a constraint on the average energy expectation value. In addition, they considered the maximum $S_{P}$ distribution for a two-level system subject to a constraint to the Gibbs state.[10, 11, 12, 13, 9] Generalizations to $N$-level systems were made soon after for the energy-constrained distribution,[8, 34] and more recently for the Gibbs state-constrained distribution.[4] [^1] We catch a minor error in the previously published GCE distribution by providing an alternative proof in section S1.2 of the SI.
Here, we critically evaluate and build upon these previous efforts with the goal of establishing a maximum entropy principle for the equilibrium wavefunction ensemble. Before proceeding, we define our criteria for equilibrium. Specifically, we will consider $P(\Gamma)$ to be a proper thermal distribution if it satisfies the three criteria originally proposed by Goldstein et al.[20] These criteria for equilibrium are:
@@
    $P(\Gamma)$ obeys the hereditary property. That is, if a total system composed of two independent subsystems (e.g. a system and a bath) has one subsystem projected out (e.g., by performing a measurement on the bath), then the projected system is time invariant, in the Gibbs state, wnd with wavefunction states distributed according to the same distribution law, $P(\Gamma)$. The hereditary property ensures that the thermal distribution is inherited by the system when measurements occur in the bath of a system-bath composite.
@@
The Scrooge ensemble is a probability distribution of pure state wavefunctions, refererenced to a density matrix, $\rho$, defined as,
$P_{\mathrm{Scr}}(\Gamma)=\frac{1}{(2\pi)^{N-1}}\frac{N!}{\det\rho}\left(\mathrm{Tr}\{\Gamma\rho^{-1}\}\right)^{-(N+1)},$ (5)
where $N$ is the deminsion of the matrices $\Gamma$ and $\rho$. This distribution was first derived by Jozsa et al. as the solution to Eq. 4 that minimizes accessible information of the system’s wavefunction, thereby being maximally ‘stingy’ with information.[27] Later, $P_{\mathrm{Scr}}(\Gamma)$ was independently derived by Goldstein et al., who demonstrated that when $\rho=\rho_{\mathrm{G}}$, $P_{\mathrm{Scr}}(\Gamma)$ uniquely satisfies the above three criteria for proper thermal equilibrium.[20] Previous work has demonstrated that the Scrooge ensemble has maximum entropy in a certain sense, as it can be obtained from a maximum entropy distribution over unnormalized quantum states (subject to a density matrix constraint), by projecting the unnormalized states to the unit sphere.[20, 31] However, the physical interpretation of this demonstration is not clear. If the Scrooge ensemble can be formulated from a maximum entropy principle on normalized wavefunctions, then the associated constraints may guide our understanding of its physical significance.
We now demonstrate that the Scrooge ensemble is the maximum entropy distribution subject to a constraint on the Rényi divergence. The Rényi divergence is a tunable measure of the difference between two probability distributions or density matrices.[39, 37] Specifically, the Rényi divergence between a pure state density matrix $\Gamma$, and an ensemble density matrix, $\rho$, is given by,
@@
where $\alpha$ is a tunable parameter capable of shifting the emphasis of the divergence between the features of $\Gamma$ and $\rho$. For $\alpha\rightarrow 0$, Eq. 6 simply measures the overlap between $\Gamma$ and $\rho$. In the limit that $\alpha\rightarrow 1$, the Rényi divergence reduces to the Kullback–Leibler divergence, or equivalently to Umegaki’s relative entropy[46] ($\lim_{\alpha\rightarrow 1}D_{\alpha}(\Gamma\|\rho)=-\mathrm{Tr}\{\Gamma\ln\rho\}$). As we increase $\alpha\rightarrow\infty$, the divergence becomes more stringent and sensitive to the largest population difference between $\Gamma$ and $\rho$.[47]
@@
where $C(\rho)$ is a yet to be specified function of $\rho$. We then determine the specific form of $C(\rho)$ that yields the equilibrium Scrooge ensemble. Using the method of Lagrange multipliers, the maximum entropy distribution is given by,
@@
where $\mu$ is the Lagrange multiplier enforcing the constraint $\langle D_{\alpha}(\Gamma\|\rho)\rangle_{P_{\alpha\mu}}=C(\rho)$. A detailed derivation of Eq. 8 appears in section S1.3 of the SI. Comparing Eq. 8 to Eq. 5 reveals that this maximum entropy solution yields the Scrooge ensemble in the case where $\alpha=2$ and $\mu=N+1$. Substituting these parameter values into $P_{\alpha\mu}(\Gamma)$, and using the resulting distribution to compute the average in Eq. 7 provides a solution to the constraint (proof in section S1.3 of the SI),
@@
In section S2 of the SI, we present a proof that the constraint in Eq. 13 is the unique constraint on the Rényi divergence that is self-consistent, establishing the Scrooge ensemble as an important maximum entropy solution among this class of constraints. That is, this particular constraint applied specifically with $\alpha=2$ is the only Rényi divergence-based constraint with a maximum entropy principle corresponding to thermal equilibrium. A consequence of this proof is that it demonstrates that any constrained Rényi divergence between the density operator and its composing wavefunctions must yield the Scrooge ensemble in a maximum entropy procedure. However, it is possible that different physical constraints exist that equivalently yield the Scrooge ensemble as a maximum entropy solution.
The constraint in Eq. 13 includes a hierarchy of indirect constraints. This includes a constraint that the first moment of $P(\Gamma)$ is equal to the Gibbs state, $\rho_{\mathrm{G}}$, which, in turn, includes an indirect constraint on the average energy of the wavefunction ensemble. In the following sections, we ascend this hierarchy of constraints and evaluate the consequences for the associated maximum entropy principles.
@@
Figure 2: The populations of the ECE (red stars) do not agree with the Gibbs state (black stars). (a) The thermal populations at high temperature are similar. (b) At low temperature, the ground state and higher excited state populations are increased in the ECE. Inset shows populations on a logarithmic scale. Results were calculated for the Hamiltonian spectrum $\sigma\{H\}=\{k\}_{k=0}^{9}$.
@@
where $\bar{H}(\Gamma)=\mathrm{Tr}\{H\Gamma\}$ is an energy associated with the state $\Gamma$, and $\beta^{\prime}$ is a statistical inverse temperature that depends on the average energy. Notably, $\beta^{\prime}$ differs from the thermodynamic inverse temperature, $\beta$. However, the two quantities can be related through the energy constraint expression (proof in section S1.1 of the SI),[22, 24, 32]
$\frac{\sum_{i}E_{i}\alpha_{i}}{\sum_{i}\alpha_{i}}+\frac{N-1}{\beta^{\prime}}=\frac{\sum_{i}E_{i}e^{-\beta E_{i}}}{\sum_{i}e^{-\beta E_{i}}},$ (15)
where
$\alpha_{i}=\frac{e^{-\beta^{\prime}E_{i}}}{\prod_{j\neq i}(\beta^{\prime}E_{j}-\beta^{\prime}E_{i})},$ (16)
and $E_{i}$ denotes the eigenvalue of the $i$th energy eigenstate.
@@
Although the Gibbs state is preserved in the GCE, we demonstrate that the GCE violates the hereditary property and thus cannot be a valid thermal distribution. The hereditary property requires the projected ensemble on the system (e.g., in a system-bath composite) to be the same as the thermal distribution on the system.[20] More concretely, let $|\Psi\rangle\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ be a wavefunction sampled from the thermal distribution $P(\Gamma)$ on the composite Hilbert space of $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$. Then the projected ensemble on $\mathcal{H}_{1}$ given a measurement basis $\{|q_{2}\rangle\}$ for $\mathcal{H}_{2}$ is
@@
Figure 3: The hereditary property is violated by the GCE and satisfied by the Scrooge ensemble. The Kolmogorov-Smirnov (KS) statistic of the projected ensemble and thermal distribution of the system as a function of inverse temperature $\beta$. The GCE satisfies the KS test at high temperatures for the significance levels of $\alpha=5\%$ and $0.5\%$, but fails at low temperatures. The Scrooge ensemble satisfies the hereditary property for all temperatures.[20] Results are calculated by Monte Carlo sampling, with the system and bath spectra given by $\sigma\{H\}=\{0,1\}$. Error bars are standard error in the mean over 1010 trials with $10^{7}$ samples.
@@ Discussion
The maximum entropy constraint of the Scrooge ensemble currently lacks a strong a priori physical justification. Due to the uniqueness of the constraint $\langle D_{2}(\Gamma\|\rho)\rangle_{P}=\langle H_{A}(\rho)\rangle_{A}$ amongst the class of divergence constraints, justifying it reduces to rationalizing why the average excess surprisal, or loosely speaking, the ‘distance’ of wavefunctions from their average $\rho$, should depend on $\rho$ in an ensemble of non-interacting and statistically independent systems at equilibrium.